√1+cosA/√1-CosA=sinA/1-cosA

Answer by amarjeeth123(574)

LHS=√1+cosA/√1-CosA
Multiplying numerator and denominator by √1-CosA we get,
LHS=√1+cosA.√1-CosA/√1-CosA.√1-CosA=√1-cosA^2/√(1-cosA)^2=√sinA^2/(1-cosA)=sinA/(1-cosA)=RHS